Solve by iteration method:
(i) x^3 - 2 x^2 − 5=0
1
Expert's answer
2020-11-11T19:07:03-0500
"x^3-2x^2-5=0"

"x=g(x)=\sqrt[3]{2x^2+5}"

"x_0=2.6"

"x_{i+1}=g(x_i)=\sqrt[3]{2x_i^2+5}, i=0, 1,2,..."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 i & x_i & g(x_i) \\ \hline
 0 & 2.6 & 2.64573897 \\
 1 & 2.64573897& 2.66839553 \\
 2 &2.66839553 & 2.67962112 \\
 3 & 2.67962112 & 2.68518353\\
 4 & 2.68518353 & 2.68793986 \\
 5 & 2.68793986 & 2.68930573\\
 6 & 2.68930573 & 2.68998257\\
 7 & 2.68998257 & 2.69031797\\
 8 & 2.69031797 & 2.69048418\\
 9 & 2.69048418 & 2.69056654\\
10 & 2.69056654 & 2.69060735\\
11 & 2.69060735 & 2.69062758\\
12 & 2.69062758 & 2.69063760\\
13 & 2.69063760 & 2.69064257\\
14 & 2.69064257 & 2.69064503\\
15 & 2.69064503 & 2.69064625\\
16 & 2.69064625 & 2.69064685\\
17 & 2.69064685 & 2.69064715\\
18 & 2.69064715 & 2.69064730\\
19 & 2.69064730 & 2.69064737\\
20 & 2.69064737 & 2.69064741\\
21 & 2.69064741 & 2.69064743\\
22 & 2.69064743 & 2.69064744\\
\hdashline
23 &2.69064744 & 2.69064744\\
\end{array}"

"x\approx2.69064744"